Fermi National Laboratory


Accelerator Update for February 28, 2005 - March 2, 2005

Monday February 28
The day shift began with the Antiproton Source (Pbar) stacking, with Tevatron (TeV) experts conducting studies, with the Recycler in proton study mode, with work in the Switchyard (SY) tunnels, and with MiniBooNE taking beam.

Operations put the TeV into shot setup at 11:16 AM.

Operations established store 4011 at 1:45 PM with an initial luminosity of 78.71E30.

Pbar, NuMI, and the Main Injector (MI) were set up to stack and send beam to NuMI at 2:26 PM. The first attempt to do this resulted in a NuMI trip. However, at 3:28 PM Operations successfully sent beam to Pbar and NuMI.

NuMI experts ended their study at 6:09 PM. Operations had Pbar switched back to normal stacking by 6:28 PM.

Operators interlocked the Meson enclosures and MCenter resumed taking beam at 8:51 PM.

Tuesday March 1
The midnight shift began with Operations monitoring store 4011, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 1:34 AM, a Booster RF station (BRF17) tripped off for the first of five times in twenty minutes. Operators left the station off after the last trip.

At 2:54 AM, BRF18 tripped off due to a vacuum burst. It reset okay.

BRF17 had been operating for a while, but then tripped again. Operators bypassed the station at 4:57 AM. RF experts had the station back on by 5:55 AM.

At 6:35 AM, a Linac RF station (LRF4) tripped off for the first of several times over twenty minutes.

MCenter experimenters asked for an access at 6:52 AM.

Pbar resumed stacking at 6:55 AM.

At 6:57 AM, BRF17 tripped off again. It reset once, but then tripped off. Operators bypassed the station.

Operators began preparing the Booster and MI for an access at 8 AM.

At 8:02 AM, operators tried to bring up BRF17, but it tripped off two more times; the station was bypassed again.

Operations turned off Booster and MI at 8:12 AM for access.

Linac experts accessed the H- Source for motor generator work at 8:44 AM. They had this work finished and the source back on within an hour.

Operations reported at 11:22 AM that all MI accesses were complete.

Booster came out of access at 11:57 AM, but a power supply problem (GMPS) delayed turn on.

At 2:37 PM, two Linac quadrupole power supplies (QPS132 & 221) tripped off. A Linac technician investigated.

At 3:22 PM, Booster techs reported that one of the GMPS power supplies, there are four in total, needed major work. They bypassed the supply.

MiniBooNE resumed taking beam at 3:48 PM and stacking resumed at 3:53 PM.

A NuMI expert began a study period at 4:17 PM and a Recycler expert began a study a minute later.

At 4:19 PM, BRF17 tripped off twice in two minutes. It reset okay, but then at 4:25 PM, it tripped off for the first of six times in thirty minutes. It was left off for a few minutes and then turned back on at 5:07 PM.

At 4:33 PM, Operations was informed that ten NuMI vacuum pumps had tripped off. The pumps were all turned back on.

At 5:14 PM, BRF17 & 18 tripped off. They both reset okay.

At 5:16 PM, a Linac RF station (KRF5) tripped off for the first of two times in fifteen minutes. It reset okay both times.

At 6:27 PM, BRF6 tripped off for the first of four times. It reset okay each time.

Operations terminated store 4011 at 7 PM and then put the TeV into shot setup.

At 9:04 PM, the TeV suffered a quench at sector B4. Operations investigated and discovered that the skew quadrupole's power supply (T:SQ) had failed. An expert was called in to make repairs.

Pbar resumed stacking at 9:29 PM.

A SY expert began tuning the beam for MCenter at 10:46 PM.

Monday March 2
The midnight shift began with power supply techs working on T:SQ, with Pbar stacking, with the Recycler conducting studies, and with MiniBooNE and MCenter taking beam.

A power supply tech reported at 12:15 AM that T:SQ was repaired.

Operations put the TeV through a dry squeeze at 12:26 AM.

At 12:50 AM, BRF6 tripped off. It reset okay.

Operations put the TeV through a wet squeeze at 1:08 AM.

A TeV expert began a study period at 1:50 AM.

At 2:11 AM, BRF7 & 8 tripped off due to a vacuum pump trip. Everything reset okay.

At 3:57 AM, LRF3 tripped off. Operators investigated and then called in an expert.

TeV studies ended at 4:44 AM.

Operations reported at 6 AM that the H- Source began having a problems. Operators investigated and then called in experts.

The Plan for Wednesday
The plan for today is to get the source working enough to get at least a small amount of beam to the TeV for study purposes and stacking. A shot setup will follow as soon as the TeV completes its study or the stack reaches 150mA.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/2/2005   email Fermilab